# 3DTopia-XL: Scaling High-quality 3D Asset Generation via Primitive Diffusion

### v1

@@ 3.1.2 Computing PrimX from Textured Mesh
We assume all textured meshes are provided in GLB format which contains triangular meshes, texture and material maps, and corresponding UV mappings. The vertices of the target mesh are first normalized within the unit cube. To initialize the position of primitives, we first apply uniform random sampling on the mesh surface to get $\hat{N}$ candidate initial points. Then, we perform farthest point sampling on this candidate point set to get $N$ valid initial positions for all primitives. This two-step initialization of position ensures good coverage of $F_{{\mathcal{V}}}$ over the boundary neighborhood ${\mathcal{U}}$ while also keeping the high-frequency shape details as much as possible. Then, we compute the L2 distance of each primitive to its nearest neighbors, taking the corresponding value as the initial scale factor for each primitive.
@@
To initialize the payload of primitives, we first compute candidate points in global coordinates using initialized positions ${\mathbf{t}}_{k}$ and scales $s_{k}$ as ${\mathbf{t}}_{k}+s_{k}{\bm{I}}$ for each primitive, where ${\bm{I}}$ is the unit local voxel grid with a resolution of $a^{3}$. To initialize the SDF value, we query the SDF function converted from the 3D shape at each candidate point, i.e., ${\bm{X}}_{k}^{\mathrm{SDF}}=F_{{\mathcal{S}}}^{\mathrm{SDF}}({\mathbf{t}}_{k}+s_{k}{\bm{I}})$. Notably, it is non-trivial to get a robust conversion from arbitrary 3D shape to volumetric SDF function. Our implementation is based on an efficient ray marching with bounding volume hierarchy that works well with non-watertight topology. To initialize the color and material values, we sample the corresponding albedo colors and material values from UV space using geometric functions $F_{{\mathcal{S}}}^{\mathrm{RGB}}$ and $F_{{\mathcal{S}}}^{\mathrm{Mat}}$. In specific, we compute the closest face and corresponding barycentric coordinates for each candidate point on 3D mesh, then interpolate the UV coordinates and sample from the texture maps to get the value.
Finetuning.
@@
Even if the initialization above offers a fairly good estimate of $F_{{\mathcal{S}}}$, a rapid finetuning process can further decrease the approximation error via gradient descent. Specifically, we optimize the well-initialized PrimX with a regression-based loss on SDF, albedo, and material values:
${\mathcal{L}}({\mathbf{x}};{\mathcal{V}})=\lambda_{\mathrm{SDF}}||F_{{\mathcal{S}}}^{\mathrm{SDF}}({\mathbf{x}})-F_{{\mathcal{V}}}^{\mathrm{SDF}}({\mathbf{x}})||_{1}+\lambda(||F_{{\mathcal{S}}}^{\mathrm{RGB}}({\mathbf{x}})-F_{{\mathcal{V}}}^{\mathrm{RGB}}({\mathbf{x}})||_{1}+||F_{{\mathcal{S}}}^{\mathrm{Mat}}({\mathbf{x}})-F_{{\mathcal{V}}}^{\mathrm{Mat}}({\mathbf{x}})||_{1}),$ (7)
@@
where $\forall{\mathbf{x}}\in{\mathcal{U}}$, and $\lambda_{\mathrm{SDF}},\lambda$ are loss weights. We employ a two-stage finetuning strategy where we optimize with $\lambda_{\mathrm{SDF}}=10$ and $\lambda=0$ for the first 1k iterations and $\lambda_{\mathrm{SDF}}=0$ and $\lambda=1$ for the second 1k iterations.
@@ 3.2 Primitive Patch Compression
In this section, we introduce our patch-based compression scheme on primitives for two main purposes: 1) incorporating inter-channel correlations between geometry, color, and materials; and 2) compressing 3D primitives to latent tokens for efficient latent generative modeling.
We opt for using a variational autoencoder (Kingma, 2013) (VAE) operating on local voxel patches which compresses the payload of each primitive into latent tokens, i.e., $F_{\mathrm{ae}}:{\mathbb{R}}^{D}\xrightarrow{}{\mathbb{R}}^{d}$. Specifically, the autoencoder $F_{\mathrm{ae}}$ consists of an encoder $E$ and a decoder $D$ building with 3D convolutional layers. The encoder $F_{\mathrm{ae}}$ has a downsampling rate of $48$ that compresses the voxel payload ${\bm{X}}_{k}\in{\mathbb{R}}^{a^{3}\times 6}$ into the voxel latent $\hat{{\bm{X}}}_{k}\in{\mathbb{R}}^{(a/2)^{3}\times 1}$. We train this variational autoencoder $F_{\mathrm{ae}}=\{E,D\}$ with KL-regularized reconstruction loss:
${\mathcal{L}}_{\mathrm{ae}}({\bm{X}};E,D)=\frac{1}{N}\sum_{k=1}^{N}[||{\bm{X}}_{k}-D(E({\bm{X}}_{k}))||_{2}+\lambda_{\mathrm{kl}}{\mathcal{L}}_{\mathrm{kl}}({\bm{X}}_{k},E)],$ (8)
where $\lambda_{\mathrm{kl}}$ is the weight for KL regularization over the latent space of the autoencoder. Note that, unlike other works on 2D/3D latent diffusion models (Nam et al., 2022; Zhang et al., 2024c; Rombach et al., 2022) that perform global compression over the collection of patches or global 3D latent, the proposed method only compresses each local primitive patch independently and defers the modeling of global semantics and inter-patch correlation to the powerful diffusion model. Once the VAE is trained, we can compress the raw PrimX as ${\mathcal{V}}_{k}=\{{\mathbf{t}}_{k},s_{k},E({\bm{X}}_{k})\}$. It leads to the low-dimensional parameter space for the diffusion model as ${\mathcal{V}}\in{\mathbb{R}}^{N\times d}$, where $d=3+1+(a/2)^{3}$. In practice, this compact parameter space significantly allows allocating more model parameters given a fixed computational budget, which is the key to scaling up 3D generative models in high resolution.
@@ 3.3 Latent Primitive Diffusion
On top of PrimX (Sec. 3.1) and the corresponding VAE (Sec. 3.2), the problem of 3D object generation is then converted to learning the distribution $p({\mathcal{V}})$ over large-scale datasets. Our goal is to train a diffusion model (Ho et al., 2020) that takes as input random noise ${\mathcal{V}}^{T}$ and conditions ${\mathbf{c}}$, and predicts PrimX samples. Note that, the target space for denoising is ${\mathcal{V}}^{T}\in{\mathbb{R}}^{N\times d}$, where $d=3+1+(a/2)^{3}$.
@@
where $\mathrm{CrossAttn}({\mathbf{q}},{\mathbf{k}},{\mathbf{v}})$ denotes the cross-attention layer with query, key, and value as input. $\mathrm{SelfAttn}(\cdot)$ denotes the self-attention layer. $\mathrm{AdaLN}(\cdot,t)$ denotes adaptive layer normalization layers to inject timestep conditioned modulation to cross-attention, self-attention, and feed-forward layers. Moreover, we employ the pre-normalization scheme (Xiong et al., 2020) for training stability. For noise scheduling, we use discrete 1,000 noise steps with a cosine scheduler during training. We opt for “v-prediction” (Salimans & Ho, 2022) with Classifier-Free Guidance (CFG) (Ho & Salimans, 2022) as the training objective for better conditional generation quality and faster convergence:
${\mathcal{L}}_{\mathrm{diff}}(\Phi)=\mathbb{E}_{t\sim[1,T],{\mathcal{V}}^{0},{\mathcal{V}}^{t}}[||(\sqrt{\bar{\alpha}_{t}}{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}{\mathcal{V}}^{0})-g_{\Phi}({\mathcal{V}}^{t},t,\bar{{\mathbf{c}}}(b))||_{2}^{2}],$ (10)
@@ 4 Experiments
Figure 4: 3DTopia-XL can generate 3D assets directly from texts or single-view images.
4.1 Implementation Details
4.1.1 Data Standardization
Datasets.
The scale and quality of 3D data determine the quality and effectiveness of 3D generative models at scales. We filter out low-quality meshes, such as fragmented shapes and large-scale scenes, resulting in a refined collection of 256k objects from Objaverse (Deitke et al., 2023b). Computing PrimX on large-scale datasets involves two critical steps: 1) Instantiation of sampling functions $\{F_{{\mathcal{S}}}^{\mathrm{SDF}},F_{{\mathcal{S}}}^{\mathrm{RGB}},F_{{\mathcal{S}}}^{\mathrm{Mat}}\}$ from a GLB file and 2) Execution of the fitting algorithm in Sec. 3.1.2. Given the massive amount of meshes from diverse sources in Objaverse, there are challenges for properly instantiating the sampling functions in a universal way such as fragmented meshes, non-watertight shapes, and inconsistent UVs. Our standardized procedure starts with loading the GLB file as a connected graph. We filter out subcomponents that have less than 3 face adjacency which typically represent isolated planes or grounds. After that, all mesh subcomponents are globally normalized to the unit cube $[-1,1]$ given one unique global bounding box. Then, we instantiate geometric sampling functions for each mesh subcomponent for SDF, texture, and material values.
PrimX Hyperparameters.
To get a tradeoff between computational complexity and approximation error, we choose our PrimX to have $N=2048$ primitives where each primitive’s payload has a resolution of $a=8$. It indicates that the sequence length of our primitive diffusion Transformer is also $2048$ where each token has a dimension of $d=3+1+(a/2)^{3}=68$. For the rapid finetuning stage for computing PrimX, we sample 500k points from the target mesh, where 300k points are sampled on the surface and 200k points are sampled with a standard deviation of 0.01 near the surface. The finetuning stage is run for 2k iterations with a batch size of 16k points using an Adam (Kingma & Ba, 2014) optimizer at a learning rate of $1\times 10^{-4}$.
4.1.2 Condition Signals
Conditioners.
The conditional generation formulation in Sec. 3.3 is compatible with most modalities. In this paper, we mainly explored conditional generation on two modalities, images and texts. For image-conditioned models, we leverage pretrained DINOv2 model (Oquab et al., 2023), specifically “DINOv2-ViT-B/14”, to extract visual tokens from input images and take it as the input condition ${\mathbf{c}}$. For text-conditioned models, we leverage the text encoder of the pretrained image-language model (Radford et al., 2021), namely “CLIP-ViT-L/14”, to extract language tokens from input texts.
Images.
Thanks to our high-quality representation PrimX and its capability for efficient rendering, we do not need to undergo the complex and expensive rendering process like other works (Hong et al., 2023), which renders all raw meshes into 2D images for training. Instead, we opt to use the front-view image rendered by Eq. 6 which is 1) efficient enough to compute on-the-fly, and 2) consistent with the underlying representation compared with the rendering from the raw mesh.
Text Captions.
We sample 200,000 data points from Objaverse to generate text captions. For each object, six different views are rendered against a white background. We then use GPT-4V to generate keywords based on these images, focusing on aspects such as geometry, texture, and style. While we pre-define certain keywords for each aspect, the model is also encouraged to generate more context-specific keywords. Once the keywords are obtained, GPT-4 is employed to summarize them into a single sentence, beginning with ’A 3D model of…’. These text captions are subsequently prepared as input conditions.
4.1.3 Model Details
Training.
We train the latent primitive diffusion model $g_{\Phi}$ using a Transformer-based architecture for scalability. Our final model is built with 28 layers with 16-head attentions and 1152 hidden dimensions, leading to a total number of $\sim$1B parameters. We train $g_{\Phi}$ with a batch size of 1024 using an AdamW (Loshchilov & Hutter, 2017) optimizer. The learning rate is set to $1\times 10^{-4}$ with a cosine learning rate warmup for 3k iterations. The probability for condition dropout is set to $b=0.1$. During training, we apply EMA (Exponential Moving Average) on the model’s weight with a decay of 0.9999 for better stability. The image-conditioned model is trained on 16 nodes of 8 A100 GPUs for 350k iterations, which takes around 14 days to converge. The text-conditioned model is trained on 16 nodes of 8 A100 GPUs for 200k iterations, which takes around 5 days to converge.
The 3D VAE for patch-wise primitive compression is built with 3D convolutional layers. We train the VAE on a subset of the entire dataset with 98k samples, finding it generalizes well on unseen data. The training takes 60k iterations with a batch size of 256 using an Adam (Kingma & Ba, 2014) optimizer with a learning rate of $1\times 10^{-4}$. Note that, this batch size indicates the total number of PrimX samples per iteration. As our VAE operates on each primitive independently, the actual batch size would be $N\times 256$. We set the weight for KL regularization to $\lambda_{\mathrm{kl}}=5\times 10^{-4}$. The training is distributed on 8 nodes of 8 A100 GPUs, which takes about 18 hours.
Inference.
By default, we evaluate our model with a 25-step DDIM (Song et al., 2020) sampler and CFG scale at 6. We find the optimal range of the DDIM sampling steps is 25 $\sim$ 100 while the CFG scale is $4\sim 10$. The inference can be efficiently done on a single A100 GPU within 5 seconds.
4.2 Representation Evaluation
@@
Evaluation Protocol.
We first evaluate different designs of 3D representations in the context of 3D generative modeling. Our evaluation principles focus on two aspects: 1) runtime from GLB mesh to the representation, and 2) approximation error for shape, texture, and material given a fixed computational budget. Given 30 GLB meshes randomly sampled from our training dataset, we take the average fitting time till convergence as runtime, which is measured as the wall time on an A100 GPU. For geometry quality, we evaluate the Chamfer Distance (CD) between the ground truth mesh and extracted mesh after the fitting and the Peak Signal-to-Noise Ratio (PSNR) of SDF values of 500k points sampled near the shape surface. For appearance quality, we evaluate the PSNR of RGB (albedo) and materials values of 500k points sampled near the surface.
Baselines.
Given our final hyperparameters of PrimX, where $N=2048,a=8$, we fix the number of parameters of all representations to $2048\times 8^{3}\approx 1.05\mathrm{M}$ for comparisons. We compare four alternative representations: 1) MLP: a pure Multi-Layer Perceptron with 3 layers and 1024 hidden dimensions; 2) MLP w/ PE: the MLP baseline with Positional Encoding (PE) (Mildenhall et al., 2020) to the input coordinates; 3) Triplane (Chan et al., 2022): three orthogonal 2D planes with a resolution of $128\times 128$ and 16 channels, followed by a two-layer MLP decoder with 512 hidden dimensions. 4) Dense Voxels: a dense 3D voxel with a resolution of $100\times 100\times 100$. All methods are trained with the same objectives (Eq. 7) and points sampling strategy (Sec. 4.1.1) as ours.
Results.
Quantitative results are presented in Table 1, which shows that PrimX achieves the least approximation error among all methods, especially for geometry (indicated by CD). Besides the best quality, the proposed representation demonstrates significant efficiency in terms of runtime with nearly 7 times faster convergence speed compared with the second best, making it scalable on large-scale datasets. Figure 5 shows qualitative comparisons. MLP-based implicit methods appear to have periodic artifacts, especially for the geometry. Triplane and dense voxels yield bumpy surfaces as well as grid artifacts around the shape surface. Instead, PrimX produces the best quality with smooth geometry and fine-grained details like the thin and tapering beard.
Figure 6: Image-to-3D comparisons. For each method, we take the textured mesh predicted from the input image into Blender and render it with the target environment map. We compare our single-view conditioned model with sparse-view reconstruction models and image-conditioned diffusion models. 3DTopia-XL achieves the best visual and geometry quality among all methods. Thanks to our capability to generate spatially varied PBR assets shown on the rightmost, our generated mesh can also produce vivid reflectance with specular highlights and glossiness.

### v2

@@ 3.1.2 Computing PrimX from Textured Mesh
Initialization. We assume all textured meshes are provided in GLB format which contains triangular meshes, texture and material maps, and corresponding UV mappings. The vertices of the target mesh are first normalized within the unit cube. To initialize the position of primitives, we first apply uniform random sampling on the mesh surface to get $\hat{N}$ candidate initial points. Then, we perform farthest point sampling on this candidate point set to get $N$ valid initial positions for all primitives. This two-step initialization of position ensures good coverage of $F_{{\mathcal{V}}}$ over the boundary neighborhood ${\mathcal{U}}$ while also keeping the high-frequency shape details as much as possible. Then, we compute the L2 distance of each primitive to its nearest neighbors, taking the value as the initial scale factor for each primitive.
@@
To initialize the payload of primitives, we first compute candidate points in global coordinates using initialized positions ${\mathbf{t}}_{k}$ and scales $s_{k}$ as ${\mathbf{t}}_{k}+s_{k}{\bm{I}}$ for each primitive, where ${\bm{I}}$ is the unit local voxel grid with a resolution of $a^{3}$. To initialize the SDF value, we query the SDF function converted from the 3D shape at each candidate point, i.e. ${\bm{X}}_{k}^{\mathrm{SDF}}=F_{{\mathcal{S}}}^{\mathrm{SDF}}({\mathbf{t}}_{k}+s_{k}{\bm{I}})$. Notably, it is non-trivial to get a robust conversion from arbitrary 3D shape to volumetric SDF function. Our implementation is based on an efficient ray marching with bounding volume hierarchy that works well with non-watertight topology. To initialize the color and material values, we sample the corresponding albedo colors and material values from UV space using geometric functions $F_{{\mathcal{S}}}^{\mathrm{RGB}}$ and $F_{{\mathcal{S}}}^{\mathrm{Mat}}$. Specifically, we compute the closest face and corresponding barycentric coordinates for each candidate point on the mesh, then interpolate UV coordinates and sample from the texture maps to get the value.
@@
Finetuning. Even if the initialization above offers a fairly good estimate of $F_{{\mathcal{S}}}$, a rapid finetuning process can further decrease the approximation error via gradient descent. Specifically, we optimize the well-initialized PrimX with a regression-based loss on SDF, albedo, and material values:
$\begin{split}{\mathcal{L}}({\mathbf{x}};{\mathcal{V}})&=\lambda_{\mathrm{SDF}}||F_{{\mathcal{S}}}^{\mathrm{SDF}}({\mathbf{x}})-F_{{\mathcal{V}}}^{\mathrm{SDF}}({\mathbf{x}})||_{1}\\ &+\lambda(||F_{{\mathcal{S}}}^{\mathrm{RGB}}({\mathbf{x}})-F_{{\mathcal{V}}}^{\mathrm{RGB}}({\mathbf{x}})||_{1}\\ &+||F_{{\mathcal{S}}}^{\mathrm{Mat}}({\mathbf{x}})-F_{{\mathcal{V}}}^{\mathrm{Mat}}({\mathbf{x}})||_{1}),\end{split}$ (9)
@@
where $\forall{\mathbf{x}}\in{\mathcal{U}}$, and $\lambda_{\mathrm{SDF}},\lambda$ are loss weights. We employ a two-stage finetuning strategy where we optimize with $\lambda_{\mathrm{SDF}}=10$ and $\lambda=0$ for the first 1k iterations and $\lambda_{\mathrm{SDF}}=0$ and $\lambda=1$ for the second 1k iterations. More details are provided in the supplementary document.
@@ 3.2 Primitive Patch Compression
In this section, we introduce our lightweight patch-based compression on individual primitives that turns 3D primitives into latent tokens for efficient generative modeling.
We opt for using a variational autoencoder [24] (VAE) operating on local voxel patches which compresses the payload of each primitive into latent tokens, i.e. $F_{\mathrm{ae}}:{\mathbb{R}}^{D}\xrightarrow{}{\mathbb{R}}^{d}$. Specifically, the autoencoder $F_{\mathrm{ae}}$ consists of an encoder $E$ and a decoder $D$ building with 3D convolutional layers. The encoder $F_{\mathrm{ae}}$ has a downsampling rate of $48$ that compresses the voxel payload ${\bm{X}}_{k}\in{\mathbb{R}}^{a^{3}\times 6}$ into the voxel latent $\hat{{\bm{X}}}_{k}\in{\mathbb{R}}^{(a/2)^{3}\times 1}$. We train $F_{\mathrm{ae}}$ with reconstruction loss:
${\mathcal{L}}_{\mathrm{ae}}({\bm{X}};E,D)=\mathbb{E}[||{\bm{X}}_{k}-D(E({\bm{X}}_{k}))||_{2}+\lambda_{\mathrm{kl}}{\mathcal{L}}_{\mathrm{kl}}({\bm{X}}_{k},E)],$ (10)
where $\lambda_{\mathrm{kl}}$ is the weight for KL regularization over the latent space. Note that, unlike other works on 2D/3D latent diffusion models [77, 50] that perform global compression over all patches, our VAE only compresses each local primitive patch independently and defers the modeling of global semantics and inter-patch correlation to the diffusion model. Once the VAE is trained, we can compress the raw PrimX as ${\mathcal{V}}_{k}=\{{\mathbf{t}}_{k},s_{k},E({\bm{X}}_{k})\}$. It leads to a low-dimensional parameter space for the diffusion model as ${\mathcal{V}}\in{\mathbb{R}}^{N\times d}$, where $d=3+1+(a/2)^{3}$. In practice, this compact parameter space significantly allows more model parameters given a fixed computational budget, which is the key to scaling up 3D generative models in high resolution.
@@ 3.3 Latent Primitive Diffusion
On top of PrimX (Sec. 3.1) and the corresponding VAE (Sec. 3.2), the problem of 3D object generation is then converted to learning the distribution $p({\mathcal{V}})$ over large-scale datasets. Our goal is to train a diffusion model [17] that takes as input random noise ${\mathcal{V}}^{T}$ and conditions ${\mathbf{c}}$, and predicts PrimX samples. Note that, the target space for denoising is ${\mathcal{V}}^{T}\in{\mathbb{R}}^{N\times d}$, where $d=3+1+(a/2)^{3}$.
@@
where $\mathrm{CrossAttn}({\mathbf{q}},{\mathbf{k}},{\mathbf{v}})$ denotes the cross-attention layer with query, key, and value as input. $\mathrm{SelfAttn}(\cdot)$ denotes the self-attention layer. $\mathrm{AdaLN}(\cdot,t)$ denotes adaptive layer normalization layers to inject timestep conditioned modulation to cross-attention, self-attention, and feed-forward layers. Moreover, we employ the pre-normalization scheme [67] for training stability. For noise scheduling, we use discrete 1,000 noise steps with a cosine scheduler during training. We opt for “v-prediction” [51] with Classifier-Free Guidance (CFG) [16] as the training objective for better conditional generation quality and faster convergence:
$\begin{split}{\mathcal{L}}_{\mathrm{diff}}(\Phi)&=\mathbb{E}_{t\sim[1,T],{\mathcal{V}}^{0},{\mathcal{V}}^{t}}[||(\\ &\sqrt{\bar{\alpha}_{t}}{\epsilon}-\sqrt{1-\bar{\alpha}_{t}}{\mathcal{V}}^{0})-g_{\Phi}({\mathcal{V}}^{t},t,\bar{{\mathbf{c}}}(b))||_{2}^{2}],\end{split}$ (12)
@@ 4 Experiments
4.1 Representation Evaluation
Evaluation Protocol. We first evaluate different designs of 3D representations in the context of 3D generative modeling. Our evaluation principles focus on two aspects: 1) runtime from GLB mesh to the representation, and 2) approximation error for shape, texture, and material given a fixed computational budget. Given 30 GLB meshes randomly sampled from our training dataset, we take the average fitting time till convergence as runtime measured on an A100 GPU. For geometry quality, we evaluate the Chamfer Distance (CD) and the Peak Signal-to-Noise Ratio (PSNR) of SDF values between ground truth meshes and fittings. For appearance quality, we evaluate the PSNR of albedo and materials values. All metrics are computed on a set of 500k points sampled near the shape surface.
Baselines. Given our final hyperparameters of PrimX, where $N=2048,a=8$, we fix the number of parameters of all representations to $2048\times 8^{3}\approx 1.05\mathrm{M}$ for comparisons. We compare four alternative representations: 1) MLP: a pure Multi-Layer Perceptron with 3 layers and 1024 hidden dimensions; 2) MLP w/ PE: the MLP baseline with Positional Encoding (PE) [39] to the input coordinates; 3) Triplane [5]: three orthogonal 2D planes with a resolution of $128\times 128$ and 16 channels, followed by a two-layer MLP decoder with 512 hidden dimensions. 4) Dense Voxels: a dense 3D voxel with a resolution of $100\times 100\times 100$. 5) Sparse Points: 2048 latent points followed by a three-layer MLP as the decoder. All methods are trained with the same objective (Eq. 9) and points sampling strategy as ours.
Figure 5: Image-to-3D comparisons. For each method, we take the textured mesh predicted from the input image into Blender and render it with the target environment map. We compare our single-view conditioned model with sparse-view reconstruction models and image-conditioned diffusion models. 3DTopia-XL achieves the best visual and geometry quality. Thanks to our capability to generate spatially varied PBR assets shown on the rightmost, our generated mesh can also produce vivid reflectance with specular highlights and glossiness.
Results. Quantitative results are presented in Table 1, which shows that PrimX achieves the least approximation error among all methods, especially for geometry (indicated by CD). Besides the best quality, the proposed representation demonstrates significant efficiency in terms of runtime with nearly 7 times faster convergence speed compared with the second best, making it scalable on large-scale datasets. Figure 4 shows qualitative comparisons. MLP-based methods have periodic artifacts, especially for the geometry. Triplane and dense voxels yield bumpy surfaces and grid artifacts. Instead, PrimX produces the best quality with smooth geometry and fine-grained details like the tapering beard.